The ratio of men to women working for a company 7 to 5 . If there are 264 employees total, how many women work for the company?
Answer by ortarico(23)

Sum of ratio: 7 + 5 = 12
--------------
number of men: 7/12 * 264 = 154
-------------
number of women: 5/12 * 264 = 110